### v1

# TRMAC: A Time-Reversal-based MAC Protocol for Wireless Networks within Computing Packages

Abstract
As chiplet-based integration and many-core architectures become the norm in high-performance computing, on-chip wireless communication has emerged as a compelling alternative to traditional interconnects. However, scalable Medium Access Control (MAC) remains a fundamental challenge, particularly under dense traffic and limited spectral resources. This paper presents TRMAC, a novel cross-layer MAC protocol that exploits the spatial focusing capability of Time Reversal (TR) to enable multiple parallel transmissions over a shared frequency channel. By leveraging the quasi-deterministic nature of on-chip wireless channels, TRMAC pre-characterizes channel impulse responses to coordinate access using energy-based thresholds, eliminating the need for orthogonal resource allocation or centralized arbitration. Through detailed physical-layer simulation and system-level evaluation on diverse traffic, TRMAC demonstrates comparable or superior performance to existing multi-channel MAC protocols, achieving low latency, high throughput, and strong scalability across hundreds of cores. TRMAC provides a low-complexity, high-efficiency solution for future Wireless Networks-on-Chip (WNoCs), particularly in chiplet-based systems where spatial reuse and modularity are critical. With simulations we prove that TRMAC can be utilized for parallel transmissions with a single frequency channel with a similar throughput and latency as in using multiple frequency bands omitting the need for complex transceivers. This work establishes a new design direction for MAC protocols that are tightly integrated with the underlying channel physics to meet the demands of next-generation computing platforms.
keywords:
Wireless Network On-chip; Time Reversal; Parallel Signal Transmission; Multi-Channel Medium Access Control
journal: Nano Communication Networksfn1fn1footnotetext: Authors gratefully acknowledge funding from the European Commission through projects with GA 101042080 (WINC), 101099697 (QUADRATURE) and 101189474 (EWiC).
1 Introduction
Recent advances in High-Performance Computing (HPC), driven by the proliferation of CPUs, GPUs, and domain-specific accelerators, have led to the integration of many cores within a single computing package, and more recently, the disaggregation of functionality across chiplets interconnected via high-density packaging [1]. To fully harness the computational capabilities of such architectures, whether monolithic or chiplet-based, parallel processing has become indispensable, enabling the concurrent execution of tasks and efficient handling of large-scale data. However, this parallelism introduces significant challenges in inter-core and inter-chiplet communication, particularly in maintaining synchronization and ensuring coherent data sharing [2, 3, 4]. Consequently, precise and efficient communication mechanisms are critical to coordinating tasks and maximizing system performance [5, 6, 3].
Wireless Networks-on-Chip (WNoCs) have recently emerged as a promising complement to traditional wired NoC architectures, offering a path to alleviate the growing challenges of latency and power dissipation in massively multi-core and System-on-Chip (SoC)/System-in-Package (SiP) designs [5, 7]. In these systems, communication is facilitated through an interposer-based environment, where antennas integrated on each core or chiplet establish wireless links within the package. Compared to conventional wired interconnects, wireless links offer substantial advantages, including single-hop data transfers, inherent broadcasting capabilities, and dynamic reconfigurability [7, 8]. However, to be viable in high-performance environments, WNoCs must meet stringent requirements: data rates exceeding 100 Gb/s, latencies below 10 ns, energy consumption below 1 pJ/bit, and Bit Error Rates (BERs) below $10^{-10}$, ensuring performance, efficiency and reliability [9].
While the use of wireless communication within chip packages offers compelling benefits, it also presents unique challenges. The confined and reflective nature of the chip environment leads to significant electromagnetic energy dispersion, resulting in high delay spread and Inter-Symbol Interference (ISI) that degrades communication at high data rates. Furthermore, simultaneous non-orthogonal transmissions over shared bandwidths introduce severe Co-Channel Interference (CCI), particularly in parallel wireless links. To ensure robust and efficient data exchange, it is crucial to adopt Medium Access Control (MAC) or multiple access schemes that begin at the physical layer to mitigate interference and avoid collisions.
Figure 1: WNoC in an interposer package with two parallel communication links. Data packets are transmitted using TR filters, which suppress CCI at neighboring nodes. Spatial channels are coordinated via TRMAC.
Fortunately, the quasi-deterministic nature of the on-chip wireless channel, enabled by fixed package dimensions and known material properties, allows for accurate pre-characterization of the Channel Impulse Response (CIR). This predictability makes Time Reversal (TR) an attractive technique for improving communication robustness [10]. By exploiting the spatial and temporal invariance of the channel, TR focuses energy both in space and time at the intended receiver as shown in Figure 1, thereby significantly reducing ISI and enhancing signal fidelity [10, 11, 12]. TR filtering employs the time-reversed version of the CIR to act as a spatial matched filter that concentrates energy at the target receiver. In reverberant environments with rich multipath effects, the uniqueness of each CIR results in stronger signal focusing and reduced interference. Additionally, TR minimizes CCI by exploiting the orthogonality of spatial channels between different transmitter-receiver pairs. This allows multiple parallel transmissions to occur simultaneously in both time and frequency domains, provided that the number of concurrent transmissions, which we refer to the Number of Parallel Transmissions (NPT), is carefully bounded. Although TR suppresses CCI, it does not eliminate it entirely, necessitating well-defined NPT thresholds, which remain stable due to the time-invariant nature of the channel.
Despite enabling spatial multiplexing, TR does not inherently prevent transmission collisions. Moreover, TR-based communication is susceptible to the deafness problem, wherein transmitters and receivers lack awareness of ongoing transmissions due to the absence of explicit coordination, beyond the NPT constraint. To address these limitations, this work introduces TRMAC: a cross-layer, adaptive MAC protocol specifically designed to manage TR-precoded transmissions in chip-scale wireless environments.
Existing WNoC MAC protocols typically rely on resource-sharing schemes such as frequency-division, time-division, or code-division multiple access (FDMA, TDMA, or CDMA) to mitigate interference in multi-link environments [13, 14]. Some approaches adopt hierarchical hybrid wired-wireless architectures, but tend to emphasize overall system design while overlooking the effects of wireless channel impairments [15, 16, 17, 18, 4]. Table 1 summarizes recent MAC protocols in terms of architecture, multiplexing, scalability, and traffic awareness capabilities. Many of these schemes are tailored to specific architectures or applications, and do not scale well with increasing core counts or dynamic traffic. Additionally, they often incur extra overhead in area and power due to complex transceiver designs or control mechanisms, and may introduce latency through intricate scheduling or arbitration.
Table 1: Summary of the design traits of existing protocols
Work Architecture Multiplexing Scalability Traffic Awareness
Shamim et al. [4] Non-static N/A High Low
Mansoor et al. [19] Non-static TDMA Medium High
Franques et al. [20] Non-static N/A High High
Mestres et al. [21] Non-static N/A High Low
Vijayakumaran et al. [13] Static [22] CDMA High High
Matolak et al. [14] Static FDMA, TDMA High Medium
DiTomaso et al. [17] Static TDMA High High
TRMAC (this work) Non-Static Spatial High High
Despite significant progress, none of the existing MAC protocols account for communication over multiple spatial channels as enabled by TR. Most designs either assume global knowledge of ongoing transmissions or rely on orthogonal resource allocation to support parallelism. In contrast, TRMAC leverages the spatial diversity offered by TR to support simultaneous transmissions in both time and frequency. It adopts a cross-layer approach, spanning physical and link layers, tailored to the unique characteristics of quasi-deterministic on-chip wireless channels. TRMAC improves system efficiency by scaling the aggregate bandwidth with the number of spatial channels, while reducing latency to meet system-level performance demands. It also complements orthogonal multiplexing methods by offering an additional degree of freedom in spectral sharing for massively multi-core and multi-chiplet systems.
In previous work [23], we analyzed the feasibility of TRMAC with respect to parallel concurrent transmissions, where the protocol can be employed with a single frequency band and multiple parallel transmissions by reducing the area and power cost on utilizing multiple frequency bands. In this work we comprehensively analyzed TRMAC with respect to the physical layer characterization, wireless data transfer and heterogeneous traffic conditions. The main contributions of this work are summarized as follows:
  • 1.
    We propose TRMAC, a cross-layer MAC protocol in detail that leverages TR filtering to enable parallel, low-interference wireless transmissions in chip-scale environments.
  • 2.
    We analytically characterize the NPT that can be supported with TR filtering under a target BER, establishing design-time thresholds based on physical-layer channel properties.
  • 3.
    We evaluate TRMAC under diverse traffic scenarios, including even, uneven, and bursty distributions, across 64 to 256 cores, demonstrating its scalability and robustness in dense multi-core and multi-chiplet systems.
The rest of the paper is organized as follows: Section 2 provides background and motivation for wireless on-chip communication and TR-based design. Section 3 details the proposed TRMAC protocol and its physical-layer implementation. Section 5 presents the performance evaluation of TRMAC against existing approaches. Section 6 reviews relevant MAC protocols in the context of WNoC architectures. Finally, Section 7 summarizes the contributions and outlines directions for future work.
2 Background
2.1 Wireless Networks-on-Chip (WNoC)
WNoC, augment traditional wired NoC infrastructures by overlaying wireless links, offering benefits such as single-hop communication, efficient broadcasting, and reconfigurability. These features are particularly valuable in large-scale systems, where dense routing and increased hop counts can lead to significant latency and power overheads. Moreover, with the rise of chiplet-based integration, where multiple heterogeneous dies are interconnected within the same package, the inter-chip communication links are less efficient than the intra-chip links, for which wireless links offer a low-latency and scalable alternative to traditional interconnect fabrics. The monolithic or chiplet-based nature of on-package systems enables controlled ElectroMagnetic (EM) propagation through well-characterized materials and structures, making it possible to tailor channel behavior using pre-defined CIRs. As a result, WNoC architectures have seen the integration of various co-designed antennas and transceivers optimized for both core-to-core and chiplet-to-chiplet wireless communication [24, 25, 26, 27, 28, 29].
To meet high-performance requirements, such as ultra-low latency and high throughput, WNoCs must efficiently utilize available bandwidth. However, modulation schemes are typically constrained by power and area, limiting practical deployments to simple schemes like On-Off Keying (OOK) [18]. Consequently, careful MAC is necessary to coordinate transmissions across shared resources and reduce contention. Several MAC protocols adapted from traditional wireless networking, such as token passing, CSMA, and TDMA, have been explored for WNoC environments as mentioned in Section 6 [30, 19, 31, 20, 17, 21]. While these approaches effectively manage channel access and collision avoidance, they rely on orthogonal resource allocation and often assume full awareness of ongoing transmissions. They fall short in exploiting the spatial reuse potential of the on-chip environment and generally overlook the deafness problem that arises in non-orthogonal, spatially diverse settings.
Figure 2: (a) TR communication between node A and node B with reduced CCI at nodes C and D. (b) Spatial and temporal focusing at the intended receiver (node B), with minimal interference at others.
2.2 Time Reversal (TR) for On-Chip Communications
TR is a physical-layer technique that focuses EM energy at a specific receiver by transmitting a signal filtered through the time-reversed version of the CIR. This process effectively leverages the multipath richness of reverberant environments to achieve spatial and temporal energy focusing. Given that both monolithic and chiplet-based packages exhibit quasi-deterministic layouts and high reverberation, TR is a highly promising technique to mitigate multipath-induced impairments.
TR offers two key benefits for on-chip communication: (1) suppression of ISI by temporal focusing, and (2) mitigation of CCI via spatial discrimination. Each transmitter–receiver pair has a unique CIR, resulting in low cross-correlation between links. As shown in Figure 2, a signal from node A pre-filtered with the time-reversed CIR to node B will be sharply focused at B, while nearby nodes (e.g., C and D) receive minimal interference. This spatial orthogonality enables multiple simultaneous transmissions over the same frequency band without the need for traditional orthogonalization.
2.3 Medium Access Control (MAC) Design in WNoC
In WNoCs, MAC protocols play a critical role in balancing latency, throughput, and energy efficiency. While traditional MAC strategies rely on orthogonal access to prevent interference, the spatial focusing offered by TR opens up new opportunities for non-orthogonal, spatially multiplexed communication. However, this also introduces new challenges such as managing interference from partially overlapping spatial channels and addressing the deafness problem, where nodes are unaware of ongoing transmissions due to focused, rather than broadcast, communication.
The design of effective MAC protocols for WNoC must account for the deterministic and static nature of the environment, enabling pre-characterized CIRs to guide communication decisions. By combining this channel knowledge with lightweight control mechanisms, it becomes feasible to manage contention across shared frequency and time resources, even in high-density multi-core and multi-chiplet systems. Additionally, the MAC must adapt to dynamic and uneven traffic patterns, which are typical in modern heterogeneous computing workloads. Therefore, simplicity, adaptability, and scalability are key requirements for practical on-chip MAC protocol designs.
3 TRMAC Protocol
3.1 Motivation
Efficient data sharing and synchronization are critical in multi-core and multi-chiplet systems. In massive SiP architectures, parallel transmissions can reduce communication latency, but limited bandwidth must be carefully shared across concurrent links, necessitating orthogonal resource allocation (e.g., frequency or time division). Existing MAC protocols for WNoCs typically rely on FDMA or TDMA to multiplex access, which either increases latency or divides limited bandwidth among competing transmissions. These approaches do not take advantage of the spatial diversity available in on-chip wireless environments. With the introduction of TR filtering, spatial multiplexing becomes feasible. By exploiting distinct CIRs among node pairs, multiple concurrent transmissions can occur on the same frequency and time without orthogonal separation, significantly increasing aggregate bandwidth and reducing latency.
Motivated by this opportunity, we introduce TRMAC, a cross-layer MAC protocol that leverages pre-characterized CIRs and the deterministic nature of on-chip channels to support parallel transmissions through spatial reuse. TRMAC establishes per-node energy thresholds and uses preamble-based signaling to manage collisions and mitigate deafness. Because it is grounded in physical-layer channel knowledge, TRMAC can be combined with FDMA or TDMA to further increase spatial reuse, offering a practical and scalable solution for high-density multi-core and multi-chiplet systems.
Figure 3: Illustration of TRMAC: (Left) successful parallel transmissions. (Middle left) preamble collisions at node 2. (Middle right) missed ACKs from Phase II. (Right) erroneous ACK affects parallel transmission.
Figure 4: Flowchart of the TRMAC protocol.
3.2 Protocol Working
The TRMAC protocol operates in three sequential phases: preamble transmission, ACK detection, and data transfer. These stages are illustrated in Figure 3, while the detailed working flowchart is presented in Figure 4.
3.2.1 Phase I: Preamble Transmission
When a packet is ready, the Transmitter (Tx) initiates communication by sending a preamble, filtered through the time-reversed CIR of the intended Receiver (Rx). The preamble contains the Tx node’s address. Upon receiving the preamble, the Rx decodes the address, selects the corresponding CIR, and returns a one-bit TR-based Acknowledgment (ACK). Due to spatial focusing, the ACK will reach only the intended Tx node. If the Rx erroneously decodes the preamble because of collision or interference, it may send an ACK to the wrong node, which will disregard it.
3.2.2 Phase II: Acknowledgment Detection
Following the preamble transmission, the Tx waits for the ACK from the Rx. Two collision cases are considered:
  • 1.
    Case 1: No ACK received. This occurs when the Rx is unable to decode the preamble correctly due to a collision or interference, or when the Rx simultaneously initiates its own transmission, leading to deafness.
  • 2.
    Case 2: ACK received with high energy. The Tx receives an ACK, but with energy significantly above the pre-characterized threshold, indicating a possible collision with another ACK. This is a rare but possible scenario if multiple transmitters simultaneously attempt communication on overlapping spatial channels.
In both cases, the Tx concludes that a collision has occurred and halts further transmission until recovery actions are performed.
3.2.3 Phase III: Data Transmission
If the Tx receives an ACK with energy matching the expected threshold, it proceeds to transmit the data packet to the Rx. A successful transmission is completed across all three phases.
3.3 Decisions and Assumptions
Transmitter: Initially, frequency channels ($f_{n}$) are uniformly distributed among nodes, with each channel covering a sector of $N_{\text{cores}}/f_{n}$. Transmissions are assumed to be globally synchronized. Each Tx node starts communication from Phase I and proceeds sequentially unless a collision is detected.
To detect ACKs and possible collisions, each Tx compares the received signal energy to a predefined threshold $E_{\text{th,correct}}$. If a Tx aims to communicate with multiple nodes simultaneously through symmetric spatial channels, this threshold is adjusted using pre-characterized values to reflect the expected composite energy.
Receiver: Upon receiving a preamble, the Rx uses threshold-based detection to identify the Tx. If multiple transmissions arrive simultaneously at the same Rx, the energy may superimpose, leading to incorrect preamble decoding and misidentification of the Tx. Thus the Rx may select the wrong CIR, causing a failed transmission and contributing to overall collision rate.
Additionally, if the intended Rx initiates its own transmission at the same time and frequency, self-interference prevents it from receiving the original preamble or sending an ACK, again contributing to the deafness problem.
A comprehensive illustration of the working of transmitter and receiver in the proposed TRMAC protocol is shown in Figure 4.
3.4 Collision Handling and Channel Awareness
When a collision is detected in Phase II, TRMAC employs a two-step recovery mechanism:
  • 1.
    Step 1: Exponential Backoff. The Tx enters a wait period that grows exponentially with each consecutive collision, reducing the probability of repeated interference during retries.
  • 2.
    Step 2: Tone-Based Signaling. A single-bit tone is broadcast over a dedicated, high-frequency channel (170 GHz, 1 GHz bandwidth) to signal the start of Phase III, as illustrated in Figure 5. If no tone is detected, the medium is assumed to be free. This mechanism allows nodes to coordinate transmission timing without additional interference or complex arbitration.
Figure 5: Timing diagram of tone-based coordination for collision recovery.
4 Methodology
The protocol leverages the unique characteristics of the on-chip wireless channel by starting from the physical layer. The framework of the methodology is illustrated in Figure 6. First, the chip package is designed and simulated using CST Microwave Studio [32] to extract the CIRs. These CIRs are pre-characterized and considered static, assuming a fixed chip layout and material composition. Second, MATLAB [33] simulations are used to implement a basic Amplitude Shift Keying (ASK) modulation and to calculate energy thresholds for bit detection with low BER. We then estimate the NPT that can coexist without exceeding acceptable BER limits, based on cross-interference among simultaneous links. Importantly, as discussed in Section 2.1, both the energy thresholds and NPT are fixed once the chip is designed and remain stable unless the physical environment changes. Finally, these parameters are incorporated into Multi2Sim [34] to evaluate protocol feasibility under practical operating conditions.
Figure 6: TRMAC framework: CIR pre-characterization enables energy-focused communication and decentralized collision management via preamble signaling and energy threshold detection.
Figure 7: BER vs. data rate for parallel spatial transmissions using TR.
4.1 Physical Layer Simulations
To model the on-chip wireless environment, a four-chiplet interposer package was simulated using CST Microwave Studio [32], with each core equipped with a single antenna operating at a center frequency of 140 GHz. The details of the flip-chip layers follows the setup described in [35]. The interposer is created by stacking four flip-chip structures on top of silicon interposer and copper layers. The vertical monopole antennas are placed on each chip-let in symmetrical positions such that the cylindrical metallic structure is passed through the silicon substrate and excited from the bottom with metal layer acting as the ground plane. After accurate antenna modeling, the CIR is obtained by sequential impulse excitations followed by the antennas, where we can extract the information of each link for post-processing. In MATLAB, serialized data is modulated using ASK, passed through a TR filter (complex conjugate of the CIR), and transmitted through the wireless channel. The receiver captures the signal and decodes the data by comparing received energy to pre-characterized energy thresholds [10, 11].
We extend this model to simulate multiple simultaneous transmissions (a MIMO scenario) to evaluate how BER varies with increasing transmission concurrency. For each transmitter, cross-interference from other links is analyzed. We determine the maximum number of spatial links (i.e. NPT) that can be reliably supported while maintaining BER below an acceptable threshold. In addition to interference, we analyze the influence of data rate. Higher data rates lead to tighter symbol spacing, increasing vulnerability to ISI, particularly in channels with high delay spread. The resulting impact of BER on date rate for multiple spatial links is presented in Figure 7. This highlights a critical trade-off between throughput and error in TR-based communication.
4.2 Wireless Transfer and Channel Allocation
Wireless data transfer in TRMAC can operate on a single or multiple frequency channels. This work focuses on a single frequency channel using three spatial channels enabled by TR. All transmissions are aligned to a global clock at 1 GHz.
Each packet consists of 80 bits: 15 for the preamble, 64 for data, and 1 for the tone as shown in Figure 5. This design supports addressing up to $2^{15}$ nodes. Communication is divided into 5 clock cycles: 1 for the preamble, 1 for ACK, and 3 for data, yielding a throughput of $80/5$ ns = 16 Gbps, which is well within the limits of achievable TR performance ($\sim$30 Gbps) with minimal CCI [10, 11].
The 140 GHz band (16 GHz bandwidth) is used for data transmission, while the 170 GHz band (1 GHz) carries the tone signal, as shown in Figure 8. This adds approximately 2 mW of power consumption to support a short bandwidth. The authors in [36], have implemented a OOK modulated non-coherant receiver obtaining 16Gb/s data rate at 60GHz antenna frequency, in 65nm CMOS. However, as TR filters have been explored from 60–200GHz center frequency [10, 12], the area and energy overhead will be approximately 1–5 mm2 and 2–5 pJ/bit/cm1/2 as evaluated in [37].
Figure 8: Channel allocation strategy in TRMAC: data at 140 GHz; tone signaling at 170 GHz.
5 Evaluation
We evaluate the performance of the proposed TRMAC protocol in comparison with two baseline MAC strategies: Reliable-Broadcast-Sensing MAC protocol for WNoC (BRS) and Token Passing [21, 20], under various synthetic traffic models and system configurations. All simulations were conducted using Multi2Sim [34], where TRMAC was implemented on a WNoC supporting spatial parallelism.
The protocols are compared in terms of latency and throughput as a function of the number of parallel transmissions, available antennas, cores, and traffic variations. TRMAC operates with spatially multiplexed links on a single frequency channel, while the baselines employ multi-channel configurations: BRS is a preamble-based MAC adapted for quasi-static wireless channels, and Token Passing creates a virtual ring among nodes using dedicated channels. To ensure fairness, we assign $N$ frequency channels to the baselines and evaluate TRMAC using $N$ simultaneous spatial transmissions on a shared frequency band. Simulation parameters are given in Table 2.
Table 2: Simulation parameters.
Parameter Value
Number of antennas 64 – 256 (one antenna per chip)
Number of parallel transmissions (NPT) 2 – 4
Number of frequency channels 1 – 4
Network configuration 80 bits per packet, 15-bit preamble
Synthetic traffic model Hurst exponent $H=0.5$ – 1
Standard deviation $\sigma=0.1$ – 10
MAC protocols evaluated TRMAC, BRS, Token Passing
Center frequency 140 GHz (data), 170 GHz (tone signaling)
Bandwidth 16 GHz (data), 1 GHz (tone)
Data rate 30 Gbps (TR with ASK modulation)
5.1 Number of Parallel Transmissions
This subsection compares the performance of TRMAC against BRS and Token Passing, with respect to the number of parallel transmissions. The core premise of TRMAC is its ability to support multiple parallel transmissions (NPT) within a single frequency channel. To evaluate its effectiveness, we assess latency and throughput for NPT values from 2 to 4, under a synthetic traffic model with burstiness and hotspotting ($H=1$, $\sigma=0.5$).
Figure 9 shows that TRMAC achieves comparable latency to BRS, and saturation throughput equivalent to Token Passing, despite operating on a single frequency channel. As the number of transmissions increases, latency in both TRMAC and BRS grows similarly due to increased contention. Token Passing exhibits the highest latency, which scales significantly with the ring length and added contention per token cycle.
These results confirm that TRMAC enables high throughput without the need for orthogonal frequency channels. Instead of reducing latency directly, it improves aggregate bandwidth by enabling concurrent transmissions in time and frequency. Moreover, TRMAC achieves performance parity with multi-channel protocols while avoiding the hardware overhead of multiple antennas and multi-band transceivers.
Figure 9: Latency-throughput curves for TRMAC and baselines with 64 cores using (a) two channels / NPT=2, (b) three channels / NPT=3 (c) four channels / NPT=4.
5.2 Number of Cores
We now evaluate the effect of increasing the number of cores (from 64 to 256) on protocol performance. We configure BRS and Token Passing with two frequency channels, while TRMAC uses NPT=4 on a single channel.
Figure 10 shows that as the number of cores increases while keeping per-core load fixed, total traffic increases proportionally. This leads to more collisions and an expansion of the token ring, degrading latency in Token Passing. Both TRMAC and BRS show lower latency at low injection rates, but Token Passing maintains higher throughput at high loads due to better collision avoidance.
However, in BRS and TRMAC, throughput slightly decreases with more cores due to increased contention. Despite this, TRMAC performs better than BRS at high injection rates, offering a favorable trade-off between latency and throughput as system size scales.
(a) Latency vs. packet injection rate.
(b) Throughput vs. packet injection rate.
Figure 10: TRMAC vs. BRS and Token Passing with increasing number of cores (64–256) using NPT=4/4 channels.
5.3 Traffic Sensitivity Analysis
We examine the sensitivity of TRMAC to spatial and temporal traffic variations using the synthetic traffic model. We analyze the standard deviation $\sigma$ (representing spatial distribution) and the Hurst exponent $H$ (representing temporal burstiness) across broad ranges: $\sigma=0.1$ 10 and $H=0.5$ 1. Results are compared with BRS, as both are designed for on-chip environments and exploit preamble-based access.
5.3.1 Spatial Distribution ($\sigma$)
The spatial distribution determines the traffic load imbalance among nodes. Lower $\sigma$ corresponds to more localized hotspot traffic, while higher values imply uniform traffic across nodes.
In Figure 11(a), TRMAC demonstrates reduced latency and increased throughput as NPT increases. More spatial channels allow better traffic dispersion. However, at higher $\sigma$, throughput slightly drops due to more concurrent access attempts causing contention.
Figure 11(b) shows that TRMAC retains stable performance as the number of cores increases from 64 to 256 under varying $\sigma$. Throughput and latency trends remain consistent, confirming robustness under spatially imbalanced traffic.
Figure 11(c) compares TRMAC with BRS across the full $\sigma$ range. Both protocols perform similarly under varying distribution, confirming TRMAC’s ability to handle both hotspot and evenly distributed traffic with efficient spectrum allocation.
Figure 11: Impact of spatial traffic distribution ($\sigma$) on (a) TRMAC 1 frequency channel with varying NPT, (b) TRMAC with increasing cores, one channel; NPT=3 (c) Comparison of TRMAC one frequency channel; NPT=3 and BRS three channels.
5.3.2 Temporal Burstiness ($H$)
The Hurst exponent $H$ represents temporal correlations in traffic. Higher values indicate more bursty behavior, while lower values suggest uniform injection.
In Figure 12(a), TRMAC maintains similar throughput and latency for both NPT=2 and NPT=3, regardless of $H$. Figure 12(b) shows that even as core count increases, the performance of TRMAC remains stable across different burstiness levels.
Figure 12(c) compares TRMAC with BRS under bursty traffic. Both protocols show nearly identical performance trends, validating TRMAC’s robustness against temporal injection variability.
Figure 12: Impact of traffic burstiness ($H$) on (a) TRMAC with one channel and varying NPT, (b) TRMAC with increasing cores, one channel; NPT=3 (c) Comparison of TRMAC one channel; NPT=3 and BRS three channels.
6 Related Work
6.1 Token Passing and CSMA
Several MAC protocols based on token-controlled data flow have been explored in prior work [38, 39], where the node holding the token gains exclusive access to the communication medium during designated time slots, eliminating the need for global synchronization. While token passing schemes typically offer high throughput under controlled conditions, they struggle to scale under hotspot or highly dynamic traffic patterns.
The fuzzy token MAC protocol proposed in [20] presents an adaptive variant that dynamically adjusts transmission probabilities based on traffic intensity, offering low-latency, collision-free transmissions under high load conditions. It incorporates a contention-aware mechanism that adapts the token region based on local collision statistics and load, thereby balancing performance under variable traffic.
A predictive time-slot allocation MAC protocol is introduced in [39], enabling on-demand medium access based on the anticipated bandwidth requirements. On the other hand, CSMA-based wireless NoC designs have been proposed in [40], while [21] presents a reliable CSMA-inspired MAC with Negative Acknowledgment (NACK) for collision detection, offering a middle ground between CSMA/CA and CSMA/CD.
Recent work in [31] compares the performance of multi-channel MAC protocols, including token passing and the broadcast-aware CSMA design from [21], under both statistical and random traffic injection patterns across varying node counts.
Although these protocols address latency and contention to varying degrees, their reliance on complex coordination or orthogonal channel usage limits scalability in massive multi-core or chiplet-based systems. In such contexts, lightweight MAC solutions capable of supporting concurrent transmissions on shared channels, such as TRMAC, offer greater scalability and efficiency.
6.2 Multiplexing
Multiplexing strategies have long been a foundational approach in WNoC MAC designs. Protocols based on TDMA, FDMA, and CDMA have been proposed to ensure low-contention data transmission through orthogonal resource allocation.
A scalable hybrid TDMA-FDMA architecture is presented in [16], with the potential to integrate emerging technologies such as carbon nanotube antennas. In [17], a time-division-based wired-wireless system is evaluated, demonstrating the feasibility of dynamic transceiver adaptation to achieve high data rates and low energy consumption.
CDMA-based approaches are explored in [13] and [30], offering the benefits of parallelism through code orthogonality. However, the requirement for complex transceiver designs capable of handling orthogonal spreading codes limits the practicality of CDMA in dense, high-speed environments.
As core and chiplet counts increase, the hardware complexity of strict multiplexing schemes becomes a critical bottleneck, motivating the need for alternative techniques like spatial multiplexing via TR filters, as employed in TRMAC.
6.3 Architecture
Hybrid wired-wireless NoC architectures have been extensively explored to mitigate the limitations of long-range wired interconnects by integrating low-latency wireless links. These designs often toggle between MAC mechanisms such as random access and token passing [19] or target application-specific goals such as shared-memory consistency and fine-grained synchronization in multi-core systems [41].
However, many of these protocols are either tied to specific architectural assumptions, tailored to particular application domains, or require intricate control logic to adapt to varying conditions. In contrast, TRMAC offers a generalized, adaptable MAC framework that leverages the quasi-static and reverberant properties of the on-chip wireless channel. By exploiting spatial reuse and energy focusing via TR filters, TRMAC can support scalable, low-complexity MAC operations across a wide range of system sizes and traffic conditions, making it a promising candidate for future many-core and multi-chiplet environments.
7 Conclusion
This work introduced TRMAC, a cross-layer MAC protocol designed for wireless communication in dense multi-core and chiplet-based systems. By leveraging the spatial focusing properties of TR in quasi-deterministic on-chip wireless channels, TRMAC enables multiple parallel transmissions on a shared frequency band without complex coordination or orthogonal resource allocation. The protocol uses pre-characterized channel impulse responses and energy thresholds to manage contention with low overhead. Physical-layer simulations and system-level evaluations show that TRMAC matches or exceeds the performance of existing MAC protocols such as BRS and token passing, achieving similar or better performance with 1/2, 1/3 of the spectrum even under bursty and hotspot traffic patterns. TRMAC share the spectral resources by multiplying the throughput with low latency. Its ability to scale with core count while avoiding the hardware complexity of multi-band transceivers makes it well-suited for future many-core and chiplet-integrated platforms. TRMAC represents a significant step toward exploiting spatial multiplexing in on-chip wireless networks and offers a robust, low-complexity solution for next-generation processor interconnects. Future directions include integrating TRMAC into heterogeneous systems, supporting QoS-aware traffic control, and aligning with chiplet interconnect standards for modular computing.